Solve
(2√xy -y)dx- xdy=0
1
Expert's answer
2020-11-19T16:22:05-0500

Given "(2\sqrt{x}y - y)dx - xdy =0"

Using seperation of variables method we have that:

"y(2\sqrt{x} - 1)dx = xdy"

Rearranging the equation we have that:

"\frac{(2\sqrt{x} - 1)dx}{x} = \frac{dy}{y}"

Taking the integral of both sides we have that:

"4\sqrt{x} - Inx + c = Iny"

which is the general solution